# Probing Top Quark - Electron Interactions at Future Colliders

### v1

@@ Results
$T{qe}[3,3,1,1].{\textrm{txt}}\equiv A_{NLO,T{qe}[3,3,1,1]}$
where $A_{NLO,T_{\phi D}}$ contains the chirality amplitudes proportional to $C_{qe}^{3311}$ and the flags $ll,rr,rl$ and $lr$= 1,0 separate the different chiralities. (Note that the relevant coefficients are contained in the files.) Color delta for incoming quarks must be included for phenomenology. The finite parts of the 1-loop amplitudes have been obtained by subtracting infrared poles (indicated as IR in the folder).
@@
$ll$, $rr$, $rl$, $lr$: Chirality flags, for example rl refers to the chirality *right* of the incoming particles and the chirality *left* of the outgoing particles .

### v2

@@ Results
$T_{qe}[3,3,1,1].{\textrm{txt}}\equiv A_{NLO,T_{qe}[3,3,1,1]}$
where $A_{NLO,T_{qe[3,3,1,1]}}$ contains the chirality amplitudes proportional to $C_{qe}^{3311}$ and the flags $l$,$r$,rlll,$r$,rl and lrlr= 1,0 separate the different chiralities. (Note that the relevant coefficients are contained in the files.) Color delta for incoming quarks must be included for phenomenology. The finite parts of the 1-loop amplitudes have been obtained by subtracting infrared poles (indicated as IR in the folder).
@@
llll, rrrr, rlrl, lrlr: Chirality flags, for example rl refers to the chirality *right* of the incoming particles and the chirality *left* of the outgoing particles .